# How to set AdamW’s weight decay as you scale model and dataset size

### v2

@@ Appendix C Weight norm
Fig. 11 left plots the final weight magnitudes against $\lambda$, with both X and Y axes log-scaled. By comparing it with the reference line $y=0.01/x$ (gray, dashed), we can see the magnitude against $\lambda$ line shares the same slope, indicating that the weight magnitude approximately follows $\mathbb{E}[|W_{i}|]=c/\lambda$, where $c$ is a constant. Fig. 11 right plots the weight magnitudes throughout optimization, where we can see that all configurations start with the same magnitude at initializaiton, but gradually converge to our predicted magnitudes as optimization proceeds.
@@
Appendix D Additional experiment results
D.1 ViT scaling experiments on ImageNet
We additionally considered training a ViT on the 320K subset of ImageNet 32x32. In particular, we used a ViT with QK-layernorm similar to the one used in the main text, and we varied the width/size of the model by multiplying the number of the hidden dimensions and internal linear layers’ width with a factor $s$. Similar to the setting in the main text, we tested the width factor in $\{0.5,1.0,2.0\}$, used a fixed $\lambda_{\text{base}}=1.0$, and swept $\eta_{\text{base}}$ in $(2.5\times 10^{-6})\times 2^{i}$ with $i\in\{0,1,\dotsc,11\}$.
@@ D.1 ViT scaling experiments on ImageNet
Figure 14: AdamW breaks the learning rate scaling of µP on ViT. Similar to the setting in Fig. 5, here we trained a ViT with different width factors on the 320K subset of ImageNet 32x32 under the direct µP scaling (Eq. 13; top row) and our proposed scaling (Eq. 15; bottom row). The direct scaling breaks the transferability of optimal $\eta_{\text{base}}$ due to changing the timescale, whereas our scaling allows optimal stable $\eta_{\text{base}}$ across model sizes.
@@
Appendix E Note on muP library
@@
Appendix F Extended experiment setups
F.1 Model specification
@@
For ViT, we adopted the implementation from https://github.com/omihub777/ViT-CIFAR/tree/main. We also incorporated QK layernorm suggested by Dehghani et al. (2023) and Wortsman et al. (2024) in order to stabilize the training when swiping learning rates. The loss function is again chosen as cross-entropy loss. Additionally, when training on CIFAR-10, we follow the suggestions in the original codebase to use auto augmentation (Cubuk et al., 2019) and label smoothing (Müller et al., 2019) with $\alpha=0.1$ to alleviate overfitting. We indeed found these techniques crucial for reaching the level of test accuracy reported by the repo. For ImageNet experiments, we used label smoothing with standard data augmentation: random cropping and horizontal flip.
@@
For width transfer experiments in Fig. 4, we used the codebase from the recent project nanoGPT-mup [^8] https://github.com/EleutherAI/nanoGPT-mup. In particular, we trained transformers with 8 layers, with a fixed attention head dimension of 64, and varied the number of attention heads between $\{4,8,16\}$. We additional included QK layernorm for better stability when swiping hyperparameters. The model is trained with a micro-batch size of 8 sequences, 16 gradient accumulation steps, and a fixed context length of 256 tokens.
F.2 Hyperparameter range

### v3

@@ Appendix D Weight norm
Fig. 11 left plots the final weight magnitudes against $\lambda$, with both X and Y axes log-scaled. By comparing it with the reference line $y=0.01/x$ (gray, dashed), we can see the magnitude against $\lambda$ line shares the same slope, indicating that the weight magnitude approximately follows $\mathbb{E}[|W_{i}|]=c/\lambda$, where $c$ is a constant. Fig. 11 right plots the weight magnitudes throughout optimization, where we can see that all configurations start with the same magnitude at initialization, but gradually converge to our predicted magnitudes as optimization proceeds.
@@
Appendix E Additional experiment results
E.1 ViT scaling experiments on ImageNet
We additionally considered training a ViT on the 320K subset of ImageNet 32x32. In particular, we used a ViT with QK-layernorm similar to the one used in the main text, and we varied the width/size of the model by multiplying the number of hidden dimensions and internal linear layers’ width with a factor $s$. Similar to the setting in the main text, we tested the width factor in $\{0.5,1.0,2.0\}$, used a fixed $\lambda_{\text{base}}=1.0$, and swept $\eta_{\text{base}}$ in $(2.5\times 10^{-6})\times 2^{i}$ with $i\in\{0,1,\dotsc,11\}$.
@@ E.1 ViT scaling experiments on ImageNet
Figure 12: Under the suggested weight decay scaling, the optimal learning rate is stable across training length. Mirroring Fig. 1 in the main text, we trained the model for 100 epochs with different dataset sizes under a fixed batch size. Using a fixed weight decay (top rows in subfig. AB), the optimal learning rate decreases with the dataset size. Under our suggested weight decay scaling (bottom rows in subfig. AB), where $\lambda\propto\frac{1}{\textrm{dataset size}}$, the optimal learning rate becomes more stable across dataset sizes. Note that we select the values for $\lambda$ as $0.1$ as they were close-to-optimal for the experiments in Fig. 1.
@@
Figure 15: AdamW breaks the learning rate scaling of µP on ViT. Similar to the setting in Fig. 5, here we trained a ViT with different width factors on the 320K subset of ImageNet 32x32 under the direct µP scaling (Eq. 13; top row) and our proposed scaling (Eq. 15; bottom row). The direct scaling breaks the transferability of optimal $\eta_{\text{base}}$ due to changing the timescale, whereas our scaling allows stable optimal $\eta_{\text{base}}$ across model sizes.
@@
Appendix F Note on muP library
@@
Appendix G Extended experiment setups
G.1 Model specification
@@
For ViT, we adopted the implementation from https://github.com/omihub777/ViT-CIFAR/tree/main. We also incorporated QK layernorm suggested by Dehghani et al. (2023) and Wortsman et al. (2024) in order to stabilize the training when sweeping learning rates. The loss function is again chosen as cross-entropy loss. Additionally, when training on CIFAR-10, we follow the suggestions in the original codebase to use auto augmentation (Cubuk et al., 2019) and label smoothing (Müller et al., 2019) with $\alpha=0.1$ to alleviate overfitting. We indeed found these techniques crucial for reaching the level of test accuracy reported by the repo. For ImageNet experiments, we used label smoothing with standard data augmentation: random cropping and horizontal flip.
@@
For width transfer experiments in Fig. 4, we used the codebase from the recent project nanoGPT-mup [^9] https://github.com/EleutherAI/nanoGPT-mup. In particular, we trained transformers with 8 layers, with a fixed attention head dimension of 64, and varied the number of attention heads between $\{4,8,16\}$. We additional included QK layernorm for better stability when sweeping hyperparameters. The model is trained with a micro-batch size of 8 sequences, 16 gradient accumulation steps, and a fixed context length of 256 tokens.
G.2 Hyperparameter range